# Grounded in Literature

### v1

@@ D.3 LLM Metrics
To validate the reliability of the LLM judgments, we conduct a small scale study with 33 report pairs and ask 6 Ph.D. students in artificial intelligence, neuroscience, and chemistry - who are unfamiliar with our system and would thus not be able to identify it in the pair - to choose the better report based on the three criteria. We provide our annotators with the same instructions given to the LLM judge. Each report pair was rated by 1, 2, or 3 annotators, and we set the final human label for each report pair based on majority vote. Reports that had equally split ratings were discarded, resulting in 18 remaining pairs. We compute the inter-annotator agreement between the human and LLM labels based on percent agreement and Cohen’s Kappa. The results shown in Table 6 underscore a relatively high percent agreement for all metrics in addition to a fair agreement based on Cohen’s kappa for both Evidence Support and Actionability.
@@
Additionally, we provide the human preference results based on the full sample of 33 report pairs in Table Table 7. Despite the low agreement on the depth metric from Table Table 6, these preference results showcase that ScholarEval is substantially preferred over o4-mini-deep-research across all three metrics and corroborate the LLM-judge preference results.
@@
LLM judge prompt You are an expert scientific reviewer. You will be given two research evaluation reports (Report A and Report B) and you must answer which one is better based on the following criteria: Evidence support: An evidence-supported report is one that grounds its claims in the literature and backs them with relevant citations that improve traceability and its overall trustworthiness and reliability. Which report has more evidence support? Actionability: Actionability is the extent to which the report offers varied, clear, and actionable suggestions that are likely to improve the research idea. Which report is more actionable? Depth: Depth is measured by the degree of engagement with the point being evaluated and the literature cited. A deep report discusses each point from multiple angles and references specific details about the literature it cites, rather than relying on generic statements followed by citations. Which response has greater depth? You must choose either A, B, or Tie (when both responses are equivalent). Read both responses thoroughly before judging, and give thorough rationales that show that you are a fair and unbiased judge. If you choose that one response is better, clearly rationalize why. If they are equivalent, then you should choose Tie. Report A: {prompt_report_a} Report B: {prompt_report_b} Judge the responses based on the criteria, while respecting the output format below:
‘‘‘json
  "Evidence-support-rationale": "...",
  "Evidence-support-winner": "<A, B, or Tie>",
  "Actionability-rationale": "...",
  "Actionability-winner": "<A, B, or Tie>",
  "Depth-rationale": "...",
  "Depth-winner": "<A, B, or Tie>"
@@
Metric Percent Agreement (%) Cohen’s $\kappa$
Evidence support 66.67 0.1818
Actionability 72.22 0.1667
Depth 61.11 -0.0161
Table 7: Human preference results on sample of 34 report pairs.
Metric ScholarEval o4-mini-deep-research Tie
@@
Appendix E Experimental setup details
E.1 Baselines
@@
Baselines Soundness Evaluation Prompt You are an expert research assistant. You are skilled at evaluating the soundness of all methods proposed in a research idea. Create an evaluation for every method being proposed to solve the research problem. If there are n methods there should be n evaluations. Methods can be planned system designs, experiments, human studies, analyses, ablations, etc. You should find related work for each method being proposed, such that it can provide evidence that supports or contradicts the method at hand. It is important that the evaluation you generate always ties back to the original research idea. Judge the support and contradictions as well as suggested actions for each method within the general context of the research idea. Be granular, making sure to reference specific details such as: algorithm/technique, datasets/inputs, computational resources, implementation details, and evaluation setup, and metrics used quantitative outcomes, comparisons to baselines/state-of-the-art, statistical significance dataset size, hardware specs, hyperparameters, or other domain-specific constraints that affect reproducibility It is required to put the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work. Here is the research idea to evaluate: [start research idea] {research_idea} [end research idea] The output should be formatted as follows, such that each extracted method from the research idea would be its own section title, in which there are three subsections: strengths, weaknesses, and suggestions.
<method_1>
"support": "evidence that supports the proposed method",
"contradictions": "evidence that contradicts the proposed
method",
"suggestions": "how can the proposed method be improved based on
the related work"
<method_2>
"support": "evidence that supports the proposed method",
"contradictions": "evidence that contradicts the proposed
method",
"suggestions": "how can the proposed method be improved based on
the related work"
<method_n>
"support": "evidence that supports the proposed method",
"contradictions": "evidence that contradicts the proposed
method",
"suggestions": "how can the proposed method be improved based
on the related work"
When searching for related work, restrict your data sources to research papers. Do not use webpages or any other sources. All papers that you reference must be published on or before cutoff_date. This is absolutely necessary. Double check every publication’s date before referencing it. The output format needs to be respected strictly. Do not change the format in any way. It is required to copy the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work.
@@
Baselines Contribution Evaluation Prompt You are an expert research assistant. You are skilled at evaluating the novelty and impact of a research idea. First, your task is to identify a small number of high-level contribution dimensions. Contribution dimensions should represent general categories of scientific contribution that are meaningful and comparable across research ideas, regardless of the field. These might include: methodology (e.g., proposing a new method, model, or procedure) application (e.g., applying existing methods to a new problem or domain) theoretical contribution (e.g., proving a new result, deriving a new model) data (e.g., constructing a new dataset, conducting original measurements or surveys) evaluation (e.g., designing an experimental protocol, benchmarking a technique) tool or system design (e.g., building software, devices, or infrastructure to support research) conceptual framework (e.g., introducing a new taxonomy or way of thinking about a problem) Do not limit your output to the examples above but rather generate suitable dimensions for the research idea given to you. Only include dimensions that are actually reflected in the research idea — do not add generic or speculative categories. Please do not generate redundant dimensions. After having identified the dimensions, search through related literature for similar work and compare the novelty and impact of the research idea to these works along each dimension. Please summarize the results of your research into separate sections with biggest strengths, weaknesses, and suggestions for improvement according to the novelty and impact of the extracted dimensions compared to related work. It is required to put the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work. Here is the research idea to evaluate: [start research idea] {research_idea} [end research idea]
<dimension_name_1>
"strengths": "direct comparison with literature where the
research idea is MORE novel and/or impactful under this
dimension",
"weaknesses": "direct comparison with literature where the
research idea is LESS novel and/or impactful under this
dimension",
"suggestions": "actionable and useful suggestions to improve
the novelty and impact of the work, if needed, based on the
evidence from the strengths and weaknesses sections"
<dimension_name_2>
"strengths": "direct comparison with literature where the
research idea is MORE novel and/or impactful under this
dimension",
"weaknesses": "direct comparison with literature where the
research idea is LESS novel and/or impactful under this
dimension",
"suggestions": "actionable and useful suggestions to improve
the novelty and impact of the work, if needed, based on the
evidence from the strengths and weaknesses sections"
<dimension_name_n>
"strengths": "direct comparison with literature where the
research idea is MORE novel and/or impactful under this
dimension",
"weaknesses": "direct comparison with literature where the
research idea is LESS novel and/or impactful under this
dimension",
"suggestions": "actionable and useful suggestions to improve
the novelty and impact of the work, if needed, based on the
evidence from the strengths and weaknesses sections"
Guidelines: When searching for related work, restrict your data sources to research papers. Do not use webpages or any other sources. All papers that you reference must be published on or before {cutoff_date}. This is absolutely necessary. Double check every publication’s date before referencing it. The output format needs to be respected strictly. Do not change the format in any way. It is required to copy the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work. Links need to be present.
E.2 ScholarEval
@@
E.3 Ablations
In our ablation experiments, we use the same setup and parameters described in subsection E.2.
@@ E.3 Ablations
Appendix F Evaluation results
F.1 Additional Results and Significance Analysis
In this section, we include detailed evaluation results supplementing the results we presented in the main paper. Namely, Table 8, 9, 10, 11, and 12 show the pairwise significance analysis of the coverage results between ScholarEval variants and every baseline overall and for each discipline. Additionally, Table 13 shows the coverage results across all systems by type, axis, severity. Table 14 outlines the latency and cost of each of ScholarEval and baseline systems.
@@ F.1 Additional Results and Significance Analysis
ScholarEval Baseline MeanSE (SD) MeanBL (SD) $p$ Sig.?
ScholarEval $_{\text{Llama-70B}}$ Llama-3.3-70B 2.04 (1.16) 1.83 (0.89) 2.6e-06 Yes
ScholarEval $_{\text{Llama-70B}}$ GPT-4.1 2.04 (1.16) 2.18 (1.08) 0.004 No
ScholarEval $_{\text{Llama-70B}}$ Claude-4-Sonnet 2.04 (1.16) 2.18 (1.04) 0.003 No
ScholarEval $_{\text{Llama-70B}}$ o4-mini-deep-research 2.04 (1.16) 2.28 (1.07) 6.6e-07 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.72 (1.47) 1.83 (0.89) 4.4e-60 Yes
ScholarEval $_{\text{GPT-4.1}}$ GPT-4.1 2.72 (1.47) 2.18 (1.08) 8.3e-22 Yes
ScholarEval $_{\text{GPT-4.1}}$ Claude-4-Sonnet 2.72 (1.47) 2.18 (1.04) 2.6e-22 Yes
ScholarEval $_{\text{GPT-4.1}}$ o4-mini-deep-research 2.72 (1.47) 2.28 (1.07) 3.4e-15 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Llama-3.3-70B 2.77 (1.40) 1.83 (0.89) 9.5e-71 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ GPT-4.1 2.77 (1.40) 2.18 (1.08) 4.0e-27 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Claude-4-Sonnet 2.77 (1.40) 2.18 (1.04) 8.4e-28 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ o4-mini-deep-research 2.77 (1.40) 2.28 (1.07) 1.7e-19 Yes
@@
ScholarEval Baseline MeanSE (SD) MeanBL (SD) $p$ Sig.?
ScholarEval $_{\text{Llama-70B}}$ Llama-3.3-70B 2.06 (1.14) 1.88 (0.92) 0.011 Yes
ScholarEval $_{\text{Llama-70B}}$ GPT-4.1 2.06 (1.14) 2.21 (1.12) 0.053 No
ScholarEval $_{\text{Llama-70B}}$ Claude-4-Sonnet 2.06 (1.14) 2.28 (1.02) 0.003 No
ScholarEval $_{\text{Llama-70B}}$ o4-mini-deep-research 2.06 (1.14) 2.35 (1.07) 1.4e-04 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.84 (1.39) 1.88 (0.92) 7.2e-30 Yes
ScholarEval $_{\text{GPT-4.1}}$ GPT-4.1 2.84 (1.39) 2.21 (1.12) 8.1e-13 Yes
ScholarEval $_{\text{GPT-4.1}}$ Claude-4-Sonnet 2.84 (1.39) 2.28 (1.02) 4.1e-11 Yes
ScholarEval $_{\text{GPT-4.1}}$ o4-mini-deep-research 2.84 (1.39) 2.35 (1.07) 1.2e-08 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Llama-3.3-70B 2.91 (1.34) 1.88 (0.92) 2.7e-35 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ GPT-4.1 2.91 (1.34) 2.21 (1.12) 5.7e-16 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Claude-4-Sonnet 2.91 (1.34) 2.28 (1.02) 3.7e-14 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ o4-mini-deep-research 2.91 (1.34) 2.35 (1.07) 3.2e-11 Yes
@@
ScholarEval Baseline MeanSE (SD) MeanBL (SD) $p$ Sig.?
ScholarEval $_{\text{Llama-70B}}$ Llama-3.3-70B 2.05 (1.18) 1.82 (0.86) 0.004 Yes
ScholarEval $_{\text{Llama-70B}}$ GPT-4.1 2.05 (1.18) 2.18 (1.10) 0.107 No
ScholarEval $_{\text{Llama-70B}}$ Claude-4-Sonnet 2.05 (1.18) 2.15 (1.08) 0.209 No
ScholarEval $_{\text{Llama-70B}}$ o4-mini-deep-research 2.05 (1.18) 2.25 (1.04) 0.011 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.74 (1.55) 1.82 (0.86) 2.0e-24 Yes
ScholarEval $_{\text{GPT-4.1}}$ GPT-4.1 2.74 (1.55) 2.18 (1.10) 5.7e-11 Yes
ScholarEval $_{\text{GPT-4.1}}$ Claude-4-Sonnet 2.74 (1.55) 2.15 (1.08) 8.2e-12 Yes
ScholarEval $_{\text{GPT-4.1}}$ o4-mini-deep-research 2.74 (1.55) 2.25 (1.04) 1.1e-08 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Llama-3.3-70B 2.55 (1.45) 1.82 (0.86) 1.4e-17 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ GPT-4.1 2.55 (1.45) 2.18 (1.10) 4.3e-06 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Claude-4-Sonnet 2.55 (1.45) 2.15 (1.08) 6.0e-07 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ o4-mini-deep-research 2.55 (1.45) 2.25 (1.04) 1.1e-04 Yes

### v2

@@ E.3 LLM Metrics
To validate the reliability of the LLM judgments, we conduct a small scale study with 33 report pairs and ask 6 Ph.D. students in artificial intelligence, neuroscience, and chemistry - who are unfamiliar with our system and would thus not be able to identify it in the pair - to choose the better report based on the three criteria. We provide our annotators with the same instructions given to the LLM judge. Each report pair was rated by 1, 2, or 3 annotators, and we set the final human label for each report pair based on majority vote. Reports that had equally split ratings were discarded, resulting in 18 remaining pairs. We compute the inter-annotator agreement between the human and LLM labels based on percent agreement and Gwet’s AC1, which is a suitable inter-annotator agreement metric for cases representing class imbalance. The results shown in Table 7 underscore a high percent agreement for all metrics and a moderate to substantial agreement based on Gwet’s AC1.
@@
Additionally, we provide the human preference results based on the full sample of 33 report pairs in Table 8. These preference results showcase that ScholarEval is substantially preferred over OpenAI Deep Research across all three metrics and corroborate the LLM-judge preference results.
@@
LLM judge prompt You are an expert scientific reviewer. You will be given two research evaluation reports (Report A and Report B) and you must answer which one is better based on the following criteria: Evidence support: An evidence-supported report is one that grounds its claims in the literature and backs them with relevant citations that improve traceability and its overall trustworthiness and reliability. Which report has more evidence support? Actionability: Actionability is the extent to which the report offers varied, clear, and actionable suggestions that are likely to improve the research idea. Which report is more actionable? Depth: Depth is measured by the degree of engagement with the point being evaluated and the literature cited. A deep report discusses each point from multiple angles and references specific details about the literature it cites, rather than relying on generic statements followed by citations. Which response has greater depth? You must choose either A, B, or Tie (when both responses are equivalent). Read both responses thoroughly before judging, and give thorough rationales that show that you are a fair and unbiased judge. If you choose that one response is better, clearly rationalize why. If they are equivalent, then you should choose Tie. Report A: {prompt_report_a} Report B: {prompt_report_b} Judge the responses based on the criteria, while respecting the output format below: ‘‘‘json { "Evidence-support-rationale": "...", "Evidence-support-winner": "<A, B, or Tie>", "Actionability-rationale": "...", "Actionability-winner": "<A, B, or Tie>", "Depth-rationale": "...", "Depth-winner": "<A, B, or Tie>" }
Metric Percent Agreement (%) Gwet’s AC1
Evidence support 66.67 0.4476
Actionability 72.22 0.1667
Depth 61.11 0.3883
@@
Metric ScholarEval OpenAI Deep Research Tie
@@
Table 8: Human preference results on sample of 34 report pairs.
Appendix F Experimental setup details
F.1 Baselines
@@
Baselines Soundness Evaluation Prompt You are an expert research assistant. You are skilled at evaluating the soundness of all methods proposed in a research idea. Create an evaluation for every method being proposed to solve the research problem. If there are n methods there should be n evaluations. Methods can be planned system designs, experiments, human studies, analyses, ablations, etc. You should find related work for each method being proposed, such that it can provide evidence that supports or contradicts the method at hand. It is important that the evaluation you generate always ties back to the original research idea. Judge the support and contradictions as well as suggested actions for each method within the general context of the research idea. Be granular, making sure to reference specific details such as: algorithm/technique, datasets/inputs, computational resources, implementation details, and evaluation setup, and metrics used quantitative outcomes, comparisons to baselines/state-of-the-art, statistical significance dataset size, hardware specs, hyperparameters, or other domain-specific constraints that affect reproducibility It is required to put the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work. Here is the research idea to evaluate: [start research idea] {research_idea} [end research idea] The output should be formatted as follows, such that each extracted method from the research idea would be its own section title, in which there are three subsections: strengths, weaknesses, and suggestions. <method_1> "support": "evidence that supports the proposed method", "contradictions": "evidence that contradicts the proposed method", "suggestions": "how can the proposed method be improved based on the related work" <method_2> "support": "evidence that supports the proposed method", "contradictions": "evidence that contradicts the proposed method", "suggestions": "how can the proposed method be improved based on the related work" <method_n> "support": "evidence that supports the proposed method", "contradictions": "evidence that contradicts the proposed method", "suggestions": "how can the proposed method be improved based on the related work" When searching for related work, restrict your data sources to research papers. Do not use webpages or any other sources. All papers that you reference must be published on or before cutoff_date. This is absolutely necessary. Double check every publication’s date before referencing it. The output format needs to be respected strictly. Do not change the format in any way. It is required to copy the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work.
@@
Baselines Contribution Evaluation Prompt You are an expert research assistant. You are skilled at evaluating the novelty and impact of a research idea. First, your task is to identify a small number of high-level contribution dimensions. Contribution dimensions should represent general categories of scientific contribution that are meaningful and comparable across research ideas, regardless of the field. These might include: methodology (e.g., proposing a new method, model, or procedure) application (e.g., applying existing methods to a new problem or domain) theoretical contribution (e.g., proving a new result, deriving a new model) data (e.g., constructing a new dataset, conducting original measurements or surveys) evaluation (e.g., designing an experimental protocol, benchmarking a technique) tool or system design (e.g., building software, devices, or infrastructure to support research) conceptual framework (e.g., introducing a new taxonomy or way of thinking about a problem) Do not limit your output to the examples above but rather generate suitable dimensions for the research idea given to you. Only include dimensions that are actually reflected in the research idea — do not add generic or speculative categories. Please do not generate redundant dimensions. After having identified the dimensions, search through related literature for similar work and compare the novelty and impact of the research idea to these works along each dimension. Please summarize the results of your research into separate sections with biggest strengths, weaknesses, and suggestions for improvement according to the novelty and impact of the extracted dimensions compared to related work. It is required to put the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work. Here is the research idea to evaluate: [start research idea] {research_idea} [end research idea] <dimension_name_1> "strengths": "direct comparison with literature where the research idea is MORE novel and/or impactful under this dimension", "weaknesses": "direct comparison with literature where the research idea is LESS novel and/or impactful under this dimension", "suggestions": "actionable and useful suggestions to improve the novelty and impact of the work, if needed, based on the evidence from the strengths and weaknesses sections" <dimension_name_2> "strengths": "direct comparison with literature where the research idea is MORE novel and/or impactful under this dimension", "weaknesses": "direct comparison with literature where the research idea is LESS novel and/or impactful under this dimension", "suggestions": "actionable and useful suggestions to improve the novelty and impact of the work, if needed, based on the evidence from the strengths and weaknesses sections" <dimension_name_n> "strengths": "direct comparison with literature where the research idea is MORE novel and/or impactful under this dimension", "weaknesses": "direct comparison with literature where the research idea is LESS novel and/or impactful under this dimension", "suggestions": "actionable and useful suggestions to improve the novelty and impact of the work, if needed, based on the evidence from the strengths and weaknesses sections" Guidelines: When searching for related work, restrict your data sources to research papers. Do not use webpages or any other sources. All papers that you reference must be published on or before {cutoff_date}. This is absolutely necessary. Double check every publication’s date before referencing it. The output format needs to be respected strictly. Do not change the format in any way. It is required to copy the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work. Links need to be present.
F.2 ScholarEval
@@
F.3 Ablations
In our ablation experiments, we use the same setup and parameters described in subsection F.2.
@@ F.3 Ablations
Appendix G Evaluation results
G.1 Additional Results and Significance Analysis
In this section, we include detailed evaluation results supplementing the results we presented in the main paper. Namely, Tables 9, 10, 11, 12, and 13 show the pairwise significance analysis of the coverage results between ScholarEval variants and every baseline overall and for each discipline. Additionally, Table 14 shows the coverage results across all systems by type, axis, severity. Moreover, we conduct evaluation on ScholarIdeas-AI using o4-mini (the underlying model of o4-mini-deep research) as the model backbone for ScholarEval, shown in Table 15. These results showcase that using the same underlying model, ScholarEval still achieves higher coverage of the rubrics in ScholarIdeas. We also conduct an evaluation of the related system that is closest in scope to the contribution module in ScholarEval, Idea Novelty Checker by (Shahid et al., 2025), which outputs novelty verdicts for ideas followed by short rationales (Table 16). This evaluation contextualizes our system against other scholarly work on idea evaluation and confirms our choice to compare ScholarEval against strong baseline that output detailed feedback at the same scope as ScholarEval (i.e. strong LLMs and deep research systems). Finally, table 17 outlines the latency and cost of ScholarEval and baseline systems.
[…]
ScholarEval $_{\text{Llama}}$ Llama-3.3-70B 2.04 (1.16) 1.83 (0.89) <<0.001 Yes
GPT-4.1 2.04 (1.16) 2.18 (1.08) 0.998 No
GPT-5.1 Instant 2.04 (1.16) 1.94 (0.98) 0.015 Yes
Claude-4-Sonnet 2.04 (1.16) 2.18 (1.04) 0.998 No
GPT-4o-search-preview 2.04 (1.16) 1.90 (0.98) 0.001 Yes
OpenAI Deep Research 2.04 (1.16) 2.28 (1.07) 1.000 No
DR Tulu 2.04 (1.16) 2.33 (1.14) 1.000 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.72 (1.47) 1.83 (0.89) <<0.001 Yes
GPT-4.1 2.72 (1.47) 2.18 (1.08) <<0.001 Yes
GPT-5.1 Instant 2.72 (1.47) 1.94 (0.98) <<0.001 Yes
Claude-4-Sonnet 2.72 (1.47) 2.18 (1.04) <<0.001 Yes
GPT-4o-search-preview 2.72 (1.47) 1.90 (0.98) <<0.001 Yes
OpenAI Deep Research 2.72 (1.47) 2.28 (1.07) <<0.001 Yes
DR Tulu 2.72 (1.47) 2.33 (1.14) <<0.001 Yes
ScholarEval $_{\text{GPT-5.1}}$ Llama-3.3-70B 2.17 (1.27) 1.83 (0.89) <<0.001 Yes
GPT-4.1 2.17 (1.27) 2.18 (1.08) 0.578 No
GPT-5.1 Instant 2.17 (1.27) 1.94 (0.98) <<0.001 Yes
Claude-4-Sonnet 2.17 (1.27) 2.18 (1.04) 0.579 No
GPT-4o-search-preview 2.17 (1.27) 1.90 (0.98) <<0.001 Yes
OpenAI Deep Research 2.17 (1.27) 2.28 (1.07) 0.985 No
DR Tulu 2.17 (1.27) 2.33 (1.14) 0.999 No
ScholarEval $_{\text{Claude}}$ Llama-3.3-70B 2.77 (1.40) 1.83 (0.89) <<0.001 Yes
GPT-4.1 2.77 (1.40) 2.18 (1.08) <<0.001 Yes
GPT-5.1 Instant 2.77 (1.40) 1.94 (0.98) <<0.001 Yes
Claude-4-Sonnet 2.77 (1.40) 2.18 (1.04) <<0.001 Yes
GPT-4o-search-preview 2.77 (1.40) 1.90 (0.98) <<0.001 Yes
OpenAI Deep Research 2.77 (1.40) 2.28 (1.07) <<0.001 Yes
DR Tulu 2.77 (1.40) 2.33 (1.14) <<0.001 Yes
@@ G.1 Additional Results and Significance Analysis
[…]
ScholarEval $_{\text{Llama}}$ Llama-3.3-70B 2.06 (1.14) 1.88 (0.92) 0.006 Yes
GPT-4.1 2.06 (1.14) 2.21 (1.12) 0.973 No
GPT-5.1 Instant 2.06 (1.14) 2.09 (0.99) 0.659 No
Claude-4-Sonnet 2.06 (1.14) 2.28 (1.02) 0.998 No
GPT-4o-search-preview 2.06 (1.14) 1.95 (0.97) 0.065 No
OpenAI Deep Research 2.06 (1.14) 2.35 (1.07) 1.000 No
DR Tulu 2.06 (1.14) 2.35 (1.11) 1.000 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.84 (1.39) 1.88 (0.92) <<0.001 Yes
GPT-4.1 2.84 (1.39) 2.21 (1.12) <<0.001 Yes
GPT-5.1 Instant 2.84 (1.39) 2.09 (0.99) <<0.001 Yes
Claude-4-Sonnet 2.84 (1.39) 2.28 (1.02) <<0.001 Yes
GPT-4o-search-preview 2.84 (1.39) 1.95 (0.97) <<0.001 Yes
OpenAI Deep Research 2.84 (1.39) 2.35 (1.07) <<0.001 Yes
DR Tulu 2.84 (1.39) 2.35 (1.11) <<0.001 Yes
ScholarEval $_{\text{GPT-5.1}}$ Llama-3.3-70B 2.31 (1.30) 1.88 (0.92) <<0.001 Yes
GPT-4.1 2.31 (1.30) 2.21 (1.12) 0.115 No
GPT-5.1 Instant 2.31 (1.30) 2.09 (0.99) 0.003 Yes
Claude-4-Sonnet 2.31 (1.30) 2.28 (1.02) 0.354 No
GPT-4o-search-preview 2.31 (1.30) 1.95 (0.97) <<0.001 Yes
OpenAI Deep Research 2.31 (1.30) 2.35 (1.07) 0.688 No
DR Tulu 2.31 (1.30) 2.35 (1.11) 0.685 No
ScholarEval $_{\text{Claude}}$ Llama-3.3-70B 2.91 (1.34) 1.88 (0.92) <<0.001 Yes
GPT-4.1 2.91 (1.34) 2.21 (1.12) <<0.001 Yes
GPT-5.1 Instant 2.91 (1.34) 2.09 (0.99) <<0.001 Yes
Claude-4-Sonnet 2.91 (1.34) 2.28 (1.02) <<0.001 Yes
GPT-4o-search-preview 2.91 (1.34) 1.95 (0.97) <<0.001 Yes
OpenAI Deep Research 2.91 (1.34) 2.35 (1.07) <<0.001 Yes
DR Tulu 2.91 (1.34) 2.35 (1.11) <<0.001 Yes
@@
[…]
ScholarEval $_{\text{Llama}}$ Llama-3.3-70B 2.05 (1.18) 1.82 (0.86) 0.003 Yes
GPT-4.1 2.05 (1.18) 2.18 (1.10) 0.923 No
GPT-5.1 Instant 2.05 (1.18) 1.85 (0.99) 0.011 Yes
Claude-4-Sonnet 2.05 (1.18) 2.15 (1.08) 0.866 No
GPT-4o-search-preview 2.05 (1.18) 1.84 (0.96) 0.007 Yes
OpenAI Deep Research 2.05 (1.18) 2.25 (1.04) 0.988 No
DR Tulu 2.05 (1.18) 2.31 (1.17) 0.997 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.74 (1.55) 1.82 (0.86) <<0.001 Yes
GPT-4.1 2.74 (1.55) 2.18 (1.10) <<0.001 Yes
GPT-5.1 Instant 2.74 (1.55) 1.85 (0.99) <<0.001 Yes
Claude-4-Sonnet 2.74 (1.55) 2.15 (1.08) <<0.001 Yes
GPT-4o-search-preview 2.74 (1.55) 1.84 (0.96) <<0.001 Yes
OpenAI Deep Research 2.74 (1.55) 2.25 (1.04) <<0.001 Yes
DR Tulu 2.74 (1.55) 2.31 (1.17) <<0.001 Yes
ScholarEval $_{\text{GPT-5.1}}$ Llama-3.3-70B 2.12 (1.24) 1.82 (0.86) <<0.001 Yes
GPT-4.1 2.12 (1.24) 2.18 (1.10) 0.739 No
GPT-5.1 Instant 2.12 (1.24) 1.85 (0.99) 0.001 Yes
Claude-4-Sonnet 2.12 (1.24) 2.15 (1.08) 0.627 No
GPT-4o-search-preview 2.12 (1.24) 1.84 (0.96) <<0.001 Yes
OpenAI Deep Research 2.12 (1.24) 2.25 (1.04) 0.922 No
DR Tulu 2.12 (1.24) 2.31 (1.17) 0.976 No
ScholarEval $_{\text{Claude}}$ Llama-3.3-70B 2.55 (1.45) 1.82 (0.86) <<0.001 Yes
GPT-4.1 2.55 (1.45) 2.18 (1.10) <<0.001 Yes
GPT-5.1 Instant 2.55 (1.45) 1.85 (0.99) <<0.001 Yes
Claude-4-Sonnet 2.55 (1.45) 2.15 (1.08) <<0.001 Yes
GPT-4o-search-preview 2.55 (1.45) 1.84 (0.96) <<0.001 Yes
OpenAI Deep Research 2.55 (1.45) 2.25 (1.04) 0.002 Yes
DR Tulu 2.55 (1.45) 2.31 (1.17) 0.011 Yes
@@
[…]
ScholarEval $_{\text{Llama}}$ Llama-3.3-70B 2.04 (1.19) 1.81 (0.90) 0.031 Yes
GPT-4.1 2.04 (1.19) 2.14 (1.04) 0.778 No
GPT-5.1 Instant 2.04 (1.19) 1.68 (0.91) 0.002 Yes
Claude-4-Sonnet 2.04 (1.19) 2.01 (1.01) 0.408 No
GPT-4o-search-preview 2.04 (1.19) 1.86 (1.00) 0.081 No
OpenAI Deep Research 2.04 (1.19) 2.08 (1.06) 0.619 No
DR Tulu 2.04 (1.19) 2.24 (1.04) 0.937 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.61 (1.42) 1.81 (0.90) <<0.001 Yes
GPT-4.1 2.61 (1.42) 2.14 (1.04) <<0.001 Yes
GPT-5.1 Instant 2.61 (1.42) 1.68 (0.91) <<0.001 Yes
Claude-4-Sonnet 2.61 (1.42) 2.01 (1.01) <<0.001 Yes
GPT-4o-search-preview 2.61 (1.42) 1.86 (1.00) <<0.001 Yes
OpenAI Deep Research 2.61 (1.42) 2.08 (1.06) <<0.001 Yes
DR Tulu 2.61 (1.42) 2.24 (1.04) 0.006 Yes
ScholarEval $_{\text{GPT-5.1}}$ Llama-3.3-70B 1.96 (1.19) 1.81 (0.90) 0.112 No
GPT-4.1 1.96 (1.19) 2.14 (1.04) 0.916 No
GPT-5.1 Instant 1.96 (1.19) 1.68 (0.91) 0.012 Yes
Claude-4-Sonnet 1.96 (1.19) 2.01 (1.01) 0.651 No
GPT-4o-search-preview 1.96 (1.19) 1.86 (1.00) 0.218 No
OpenAI Deep Research 1.96 (1.19) 2.08 (1.06) 0.819 No
DR Tulu 1.96 (1.19) 2.24 (1.04) 0.984 No
ScholarEval $_{\text{Claude}}$ Llama-3.3-70B 2.64 (1.40) 1.81 (0.90) <<0.001 Yes
GPT-4.1 2.64 (1.40) 2.14 (1.04) <<0.001 Yes
GPT-5.1 Instant 2.64 (1.40) 1.68 (0.91) <<0.001 Yes
Claude-4-Sonnet 2.64 (1.40) 2.01 (1.01) <<0.001 Yes
GPT-4o-search-preview 2.64 (1.40) 1.86 (1.00) <<0.001 Yes
OpenAI Deep Research 2.64 (1.40) 2.08 (1.06) <<0.001 Yes
DR Tulu 2.64 (1.40) 2.24 (1.04) 0.003 Yes